A dinner theatre sells two types of tickets. Floor seats cost $50 while stadium seats cost $35. If the theatre sold 170 seats for $6625, how many tickets of each type were sold?
Tried:
x = # floor seats sold
170 - x = # stadium seats sold
Can't figure out where to go from here. Help!

Answer by geetha_rama(94)

Let x be the no: of floor seats
and y be the number of stadium seats
x+y = 170 -> equation 1
Cost of floor seat = $50
Cost of stadium seat = $35
theatre sold 170 tickets for $6625
50x + 35y = 6625
=> 10x + 7y = 1325
7x+7y = 1190 (multiply equation 1 by 7)
Subtract the above 2 equations,
3x = 135
=> x = 45
y = 170 -x = 125
There are 45 floor seats and 125 stadium seats